OpenStudy (anonymous):

Alexa is closing on a house on October 12. The buyer owns the property on the day of the closing. The selling price of the home is $300,200. Alexa was accepted for a 30-year fixed-rate mortgage for $278,200 at 4.75% interest. The seller has paid $3,833.62 in property taxes for the coming year. How much will Alexa owe in prorated taxes and interest? $2,982.00 $724.00 $850.50 $1,574.50

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

take your time

jimthompson5910 (jim_thompson5910):

so Alexa is the new owner? right? or is she the seller?

OpenStudy (anonymous):

the owner i think

jimthompson5910 (jim_thompson5910):

hmm ok, yeah that makes sense because she gets the mortgage ok still thinking

jimthompson5910 (jim_thompson5910):

well the owner (the previous owner) paid "$3,833.62 in property taxes for the coming year" however, the owner has only lived in the house from Jan 1 to Oct 11 of that year. The rest of the year, from Oct 12 onto Dec 31, Alexa is now living there. Agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

how many days are there from Jan 1 to Oct 11

OpenStudy (anonymous):

depends if its a leap year haha

jimthompson5910 (jim_thompson5910):

let's assume 365 days (non leap year)

jimthompson5910 (jim_thompson5910):

you can use a calendar or use wolfram alpha to get the answer to that question type in "Jan 1 to Oct 11" into wolfram

OpenStudy (anonymous):

283 days

jimthompson5910 (jim_thompson5910):

add one to that because they don't include Jan 1 itself so 284 days

jimthompson5910 (jim_thompson5910):

how many days are from Oct 12 to Dec 31

OpenStudy (anonymous):

80 days not including oct 12 so 81

jimthompson5910 (jim_thompson5910):

notice how 284+81 = 365, so that checks out

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

so the old owner owned the house for only 284 days out of 365 days 284/365 = 0.77808 the old owner has lived there about 77.8% of the year

OpenStudy (anonymous):

and she has lived there for 22.2% right?

jimthompson5910 (jim_thompson5910):

so it's only fair to have the old owner pay only 77.8% of the taxes the new owner (Alexa) would pay the remainder

jimthompson5910 (jim_thompson5910):

yes roughly

jimthompson5910 (jim_thompson5910):

how much in taxes would Alexa have to pay for that year?

OpenStudy (anonymous):

i know i probably sound stupid but i have no idea :(

OpenStudy (anonymous):

would i do the 22% of the3,833.62

OpenStudy (anonymous):

oh okay i got it so it's c?

jimthompson5910 (jim_thompson5910):

that's just taxes though, we haven't considered interest

OpenStudy (anonymous):

oh true.

jimthompson5910 (jim_thompson5910):

hmm one sec

jimthompson5910 (jim_thompson5910):

do you know how to find the monthly payment for the mortgage?

OpenStudy (anonymous):

yes is it the p=i(1+i)^nt / (1+i)^nt -1 ??

jimthompson5910 (jim_thompson5910):

that sounds right, so what monthly payment do you get

OpenStudy (anonymous):

would it be compounding annually or monthly?

jimthompson5910 (jim_thompson5910):

monthly I think

OpenStudy (anonymous):

$1451.22

jimthompson5910 (jim_thompson5910):

the mortgage has an interest rate of 4.75% divide this by 12 to get ???

OpenStudy (anonymous):

.0039583

jimthompson5910 (jim_thompson5910):

now multiply that by the monthly payment and you get ???

OpenStudy (anonymous):

5.74

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

hmm that seems so very low

jimthompson5910 (jim_thompson5910):

maybe I messed up and maybe it's 4.75% of the payment

OpenStudy (anonymous):

68.92

jimthompson5910 (jim_thompson5910):

the rest of that is the principal which pays off the mortgage so we will have a balance of 278,200 - 68.92 = 278,131.08 at the end of Oct (after Alexa makes that monthly mortgage payment)

OpenStudy (anonymous):

so do i calculate the interest of the balance?

jimthompson5910 (jim_thompson5910):

yes what is the interest payment for the month of Nov

OpenStudy (anonymous):

after we figure this out i have another question up and i have the answer to it and how i got it could you tell me if i did it right? if not its okay.

jimthompson5910 (jim_thompson5910):

sure, but let's figure this one out

OpenStudy (anonymous):

1099.82?

OpenStudy (anonymous):

i used an amortization table so that may be wrong.

jimthompson5910 (jim_thompson5910):

no, the monthly payment stays the same

jimthompson5910 (jim_thompson5910):

$1451.22 is the monthly payment throughout the entire mortgage

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

so we are calculating the monthly interest amount?

jimthompson5910 (jim_thompson5910):

hmm oh wait, I know what I did wrong

jimthompson5910 (jim_thompson5910):

you take 0.0475/12 = .0039583 multiply that with the remaining balance (which is originally the entire mortgage)

jimthompson5910 (jim_thompson5910):

sorry we'll have to back up and redo the interest amount for Oct

OpenStudy (anonymous):

thats okay! so .0039583 x 278200 =1101.21

jimthompson5910 (jim_thompson5910):

that's your interest payment for Oct

jimthompson5910 (jim_thompson5910):

the remaining balance is the $278,200 - $1,101.21 = ??

OpenStudy (anonymous):

then for november it is .0039583 x remaining balance of 277098.79= 1096.85

OpenStudy (anonymous):

do i do december too?

jimthompson5910 (jim_thompson5910):

that's the interest payment for Nov

jimthompson5910 (jim_thompson5910):

yes til the end of the year

OpenStudy (anonymous):

276001.50 x .0039583 = 1092.51

jimthompson5910 (jim_thompson5910):

ok so we know this taxes: 851.06364 Interest owed in Oct: 1101.21 Interest owed in Nov: 1096.85 Interest owed in Dec: 1092.51

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

the problem is that when we add them up, they go over the answer choices

jimthompson5910 (jim_thompson5910):

hmm not sure what went wrong

OpenStudy (anonymous):

i dunno. :/

jimthompson5910 (jim_thompson5910):

oh wait, I'm not thinking lol

jimthompson5910 (jim_thompson5910):

the interest for Nov and Dec isn't owed until those months come up

jimthompson5910 (jim_thompson5910):

but the interest for Oct is split up so we split 1101.21 up in such a way that it is fair to the buyer and seller

jimthompson5910 (jim_thompson5910):

for Oct, the old owner lived in the house from Oct 1 to Oct 11 that's 12 days (11-1+1 = 12) Alexa lives in the house from Oct 12 to Oct 31 (19 days) so 19/31 = 0.612903 multiply that by 1101.21 to get 1101.21*0.612903 = 674.93491263 then you add this to the amount she has to pay in taxes 851.06364+674.93491263 = 1,525.99855263 it's a bit off, but this makes sense to me a bit better. So I'm thinking it's $1,574.50 and we may have made round off errors, not sure

OpenStudy (anonymous):

thank you so much!!! you don't understand how much of a help you've been to me.

jimthompson5910 (jim_thompson5910):

honestly I can't say for certain but I think it's correct. I would ask the teacher to confirm

OpenStudy (anonymous):

i will. it's probably correct though.